### v1

# Robust CATE Estimation Using Novel Ensemble Methods

Abstract
The estimation of Conditional Average Treatment Effects (CATE) is crucial for understanding the heterogeneity of treatment effects in clinical trials. We evaluate the performance of common methods, including causal forests and various meta-learners, across a diverse set of scenarios revealing that each of the methods fails in one or more of the tested scenarios. Given the inherent uncertainty of the data-generating process in real-life scenarios, the robustness of a CATE estimator to various scenarios is critical for its reliability. To address this limitation of existing methods, we propose two new ensemble methods that integrate multiple estimators to enhance prediction stability and performance - Stacked X-Learner which uses the X-Learner [10] with model stacking for estimating the nuisance functions, and Consensus Based Averaging (CBA), which averages only the models with highest internal agreement. We show that these models achieve good performance across a wide range of scenarios varying in complexity, sample size and structure of the underlying-mechanism, including a biologically driven model for PD-L1 inhibition pathway for cancer treatment.
1 Introduction
Clinical trials aim to estimate the safety and efficacy of a treatment, usually in comparison to some control. Efficacy is measured by one or more clinical outcomes which are commonly referred to as the trial’s Endpoints (EP). The main measure of interest in terms of efficacy is the Treatment Effect, which is the expected difference in outcome under treatment vs outcome under control. Since for each patient only one potential outcome is observed, a natural way to infer the treatment effect is to use multiple patients - some given the treatment and some the control and compare their outcomes on average. If the assignment of treatment is independent of the potential outcomes (i.e. the probability of being treated is not affected by the expected efficacy of the treatment), then the difference between average outcome of the treated patients and the average outcome of the control patients equals the Average Treatment Effect (ATE) [2].
[…]
Where $A$ is the treatment assignment (0 for control, 1 for treatment), and $Y^{a}$ is the outcome under treatment assignment $a$. In clinical trials, the independence of treatment from potential outcomes is achieved by randomization.
Often, there is considerable variability in the treatment effect between different patients [3, 4], making the ATE an insufficient indicator for the treatment’s performance. In such cases, the Conditional Average Treatment Effect (CATE) - the expected treatment effect, conditional on the patient’s features, $X$ is of great interest.
CATE estimation has been intensively studied in recent years in other fields [5, 6, 10, 15], as well as in the clinical field [7]. Two main approaches that have gained popularity are Meta-Learners and Causal Forests. Meta-Learners decompose the problem of estimating the unobserved causal effect into sub-problems of predicting nuisance functions - the different potential outcomes, and the propensity. Meta-Learners can leverage any Machine-Learning (ML) model to estimate the nuisance functions. Several different frameworks were suggested on how to combine these different estimation to obtain a CATE estimator (including S-Learner, T-Learner and X-Learner which are all described in section 3, among others). Causal Forests generalize Random-Forests for the task of CATE estimation by adapting the evaluation and splitting criteria to consider the expected difference in outcomes between treatments instead of the expected outcome. These methods can differ quite substantially from each other and different models have been shown to outperform/under-perform compared to others in different settings [10, 6].
In the analysis of real datasets, the underlying Data-Generation-Process (DGP) is unknown. Furthermore, since CATE cannot be directly observed, standard cross-validation methods frequently used in prediction settings are not applicable. This emphasizes the importance of CATE estimation methods that perform well across various scenarios. To this end we propose two new ensemble methods for CATE estimation - Stacked X-Learner which is an application of the X-Learner Meta-Learner using stacked ML predictors for the nuisance functions, and the Consensus Based Averaging (CBA) which compares several CATE estimates, selects a subset of models that yield the estimates with highest internal agreement and averages them. Both methods are described in detail in section 3. A wider comparison of additional ensemble methods will be explored in future work.
Since our interest is in the analysis of clinical trials data, we designed our evaluation process to emphasise the unique challenges and characteristics of such datasets. In addition to using various additive models based on combinations of linear and nonlinear transformations of features, we also used a biologically driven DGP, which utilises a graphical functional model built to describe the Programmed-Death-Ligand 1 (PD-L1) inhibitor in cancer treatment. A distinct characteristic of clinical trials, especially in Phase II exploratory trials, are their relatively small sample sizes, which make it challenging to learn complex relations from data. Furthermore, it makes it difficult to utilize techniques involving sample splitting. Therefore, this paper focuses on small datasets, ranging from 100 to 750 patients, which represent the setting of interest, as opposed to existing comparative analyses of CATE that consider a higher number of observations [5, 12].
One of the main tasks in the analysis of clinical data is Subgroup identification [9] - the process of identifying a suitable target population for a treatment. We hypothesize that good CATE estimation can be beneficial for guiding the subgroup identification process. In order to keep the focus of this paper on CATE estimation, we leave an in-depth analysis of model-based subgroup search to future work, and only propose a designated evaluation method for CATE estimators, which highlights their accuracy on potential subgroups.
In the next section we introduce the relevant notation and terminology. Section 3 describes in detail the synthetic scenarios used in the analysis, which CATE estimators were compared and the metrics used for the estimators’ evaluation. Results are then presented in section 4. This work focuses on continuous endpoints. Future work will be dedicated to other common endpoint types.
2 Notation and Terminology
We Assume each patient $i$ can be described using:
  • $x_{i}$ a vector of patient features/covariates. We assume the dimension of the feature space is $P$ and denote the $p$-th feature of the $i$-th patient as $x^{p}_{i}$.
  • $a_{i}$ treatment assignment/allocation. We limit the discussion in this paper to binary treatments, $a_{i}=0$ indicates assignment of patient $i$ to the control and $a_{i}=1$ indicating assignment to treatment.
  • $y_{i}$ the patient’s measured EP value. We refer to this also as outcome/response interchangeably. $y_{i}$ is a random function of $x_{i}$ and $a_{i}$.
Adopting the potential outcomes framework, we also assume each patient has:
  • $Y^{0}_{i}$, $Y^{1}_{i}$ the patient’s hypothetical response under control or treatment, respectively. For control patients $y_{i}=y^{0}_{i}$ and for treated patients $y_{i}=y^{1}_{i}$.
  • $\tau_{i}=y^{1}_{i}-y^{0}_{i}$ the patient’s Individual Treatment Effect (ITE). This is an unobserved quantity, since we can only observe one potential outcome for each patient.
  • $\tau(x_{i})=\operatorname{\mathbb{E}}[Y^{1}-Y^{0}|X=x_{i}]$ the Conditional Average Treatment Effect based on the patients features. By definition $\tau(x)=\operatorname{\mathbb{E}}[\tau|X=x]$.
We follow the framework used by Lipkovich et al. [9] for decomposing $Y$ into prognostic and predictive components (with some changes in notation for internal consistency):
$\operatorname{\mathbb{E}}[Y|X,A]=\gamma(X)+\tau(X)\cdot A,$
with $\gamma(X)$ being the prognostic function, which describes the expected outcome under control, and $\tau(X)$ (CATE) being the predictive factor - the expected difference between outcomes under treatment vs. under control.
3 Methods
Our study compares the performance of different CATE estimation models in different scenarios. However, the treatment effect is not directly observable in real patients, since only one potential outcome can be measured per patient. Therefore, simulated data is required in order to test the performance in a valid and reliable manner.
We conduct a thorough simulation study on a diverse set of synthetic data generating models. This section reviews in detail the content of the study - the simulated scenarios, the methods tested, and the metrics used to evaluate their performance.
3.1 Data Generation Process
In order to evaluate the robustness of the models, we employed two different types of DGPs. The first type of DGP we employed utilizes mechanistic disease models [13] that form concise descriptions of some main pathway or mechanism of action for the progression of a selected disease. Such models require in depth biological knowledge and are costly to construct, but can generate more realistically distributed data sets with complex interactions formed by the mechanistic relations defined in the model. We utilized a simplified mechanistic model for the PD-L1 pathway in urothelial cancer [14]. This model describes the connections between a few key factors related to the progression of urothelial tumors and specifically to their treatment using PD-L1 inhibitors. The model is described in detail in appendix C.
The second type of DGP is based on a linear model with the addition of nonlinear transformations and second order interactions between features.
$\begin{split}Y=\sum\limits_{j}{\beta^{j}f^{j}(X^{j})}\ +\sum\limits_{j,k}{\beta^{jk}f^{j}(X^{j})f^{k}(X^{k})}\ +\\ \left({\sum\limits_{j}{\delta^{j}g^{j}(X^{j})}\ +\sum\limits_{j,k}{\delta^{jk}g^{j}(X^{j})g^{k}(X^{k})}}\right)\cdot A+\ \epsilon,\end{split}$ (1)
where $\beta^{j}$, $\beta^{jk}$, $\delta^{j}$ and $\delta^{jk}$ are constants, $f^{j}$ and $g^{j}$ nonlinear transformations of feature $X^{j}$, and $\epsilon$ a Gaussian random variable.
This approach allows simple control over the level of complexity of the signal and signal to noise ratios. Additionally, due to its linear form, its interpretation is relatively intuitive. We present the results for 6 representative scenarios based on this approach, consisting of three “types”:
  • Linear: Scenarios that include linear prognostic and predictive components. i.e. a DGP of the form:
    $Y=\sum\limits_{p\in\{1,2..P\}}{\beta^{p}X^{p}}\ +\sum\limits_{p\in\{1,2..P\}}{\delta^{p}X^{p}}\cdot A+\ \epsilon.$
  • Slightly nonlinear: These scenarios contained nonlinear transformations for the predictive component, with no second-order interactions, resulting in the following form:
    $Y=\sum\limits_{p\in\{1,2..P\}}{\beta^{p}X^{p}}\ +\sum\limits_{p\in\{1,2..P\}}{\delta^{p}g^{p}(X^{p})}\cdot A+\ \epsilon.$
  • Highly nonlinear: Scenarios employing the full complexity of the model presented in 1, allowing for non-linearities in both prognostic and predictive features, including second order interactions.
All DGPs were used to generate training sets of sizes that are representative of clinical trial datasets, ranging from 100 patients to 750 patients. The performance was then evaluated on held-out test sets of 5000 patients.
3.2 Metrics
The evaluation of the different methods’ accuracy employed two metrics, each aimed at measuring a different performance aspects:
  • sRMSE: scaled Root Mean Squared Error
    $sRMSE:=\frac{\sqrt{\sum_{i}(\hat{\tau}_{i}-\tau_{i})^{2}/n}}{\sqrt{\sum_{i}(\tau_{i}-\bar{\tau})^{2}/n}}=\sqrt{\frac{\sum_{i}(\hat{\tau}_{i}-\tau_{i})^{2}}{\sum_{i}(\tau_{i}-\bar{\tau})^{2}}}$
    This is aimed at measuring the distance between predictions and actual CATE. Similar to RMSE, larger errors contribute substantially more to the criteria compared to smaller errors. We scale the RMSE by the standard deviation of the CATE to obtain a scenario invariant metric that is easily interpretable and gives a sense of the relationship between the typical estimation error and the overall dynamic range of the predicted values.
  • RoD: Rate of Discordance
    $RoD:=\frac{1-K(\hat{\tau},\tau)}{2}$
    Where $K$ marks Kendall’s rank correlation coefficient, usually referred to as $\tau_{Kendall}$ (we replace with $K$ here to avoid confusion with the CATE). We call this the rate of discordance because it is equivalent to the proportion of pairs (out of all possible pairs of observations) that are incorrectly ordered by the estimator.
In addition to calculating these metrics over individuals, we also propose an adaptation that averages errors over potential subgroups (SG). We define a set of subgroups of interest (for example - all half-open boxes defined by two variables and having some minimal size) and then compose all such subgroups and treat each subgroup as an effective “observation”, using the following transformations:
$\tau_{i}\mapsto\tau_{SG}:=\frac{\sum_{i\in{SG}}\tau_{i}}{|SG|}$
and
$\hat{\tau}_{i}\mapsto\hat{\tau}_{SG}:=\frac{\sum_{i\in{SG}}y_{i}\cdot a_{i}}{\sum_{i\in{SG}}a_{i}}-\frac{\sum_{i\in{SG}}y_{i}\cdot(1-a_{i})}{\sum_{i\in{SG}}(1-a_{i})}.$ (2)
We then calculate the same accuracy metrics (sRMSE and RoD) using $\tau_{SG}$ and $\hat{\tau}_{SG}$ as the actual and predicted values, accordingly. For subgroups, a natural CATE is observable, unlike the CATE of single observation, which cannot be measured in real life applications. These metrics are more relevant for evaluating performance on a predetermined set of subgroups. CATE estimation methods which perform well on the subgroup metrics can be recommended when the aim of the CATE estimation is to serve as a basis for identifying subgroups.
3.2.1 Subgroup Definition
The set of subgroups we defined as a basis for this analysis are all 2-dimensional half-open boxes in the feature space, defined by filtering on the $\sqrt{20\%}$-th quantile of one feature to obtain a subset $S_{1}$, and sequentially on the $\sqrt{20\%}$-th quantile of another feature (within $S_{1}$). This yields (when the features are independent) SGs of 20% of the population which are then evaluated using the different accuracy metrics.
3.3 Estimators
This analysis uses two general types of estimators - Causal Forests [15] and Meta-Learners [10]. For each, we have explored a wide range of possible configurations with the aim of reaching robust performance across scenarios.
Causal Forest (CF) is an adaptation of random forests - an average of regression trees. In causal forest, the base trees composing the forest are causal trees, aimed at estimating local differences between average potential outcomes [11]. Adapting ideas of doubly robust estimation for CATE [16], CF also has debiased implementations for mitigating potential bias caused by non-random treatment allocation [17].
Meta-Learners are an estimation framework for treatment effect, that enables using any ML model as a “base learner” for learning various nuisance functions and composing an estimator for CATE using a transformation of the learned functions [10, 19]. There are several common meta-learner structures, out of which we compared the following:
  • S learner: A single model is trained to regress the outcomes on the features and the treatment assignment (the treatment is treated as an additional binary variable attached to $X$).
    $\hat{y}(x,a)=\hat{\operatorname{\mathbb{E}}}[Y(X,A)|X=x,A=a]$
    Then, ATE is estimated by contrasting this model’s predictions for both potential outcomes:
    $\hat{\tau}(x)=\hat{y}(x,1)-\hat{y}(x,0)$
  • T learner: This approach uses base-learners to estimate the conditional expectations of the two potential outcomes separately - $\{(X,Y)\ ;\ A=0\}$ are used to train $\hat{\mu}_{0}(X)$, an estimator for $\operatorname{\mathbb{E}}[Y|A=0,X]$ and $\{(X,Y)\ ;\ A=1\}$ to train $\hat{\mu}_{1}(X)$ an estimator for $\operatorname{\mathbb{E}}[Y|A=1,X]$. Finally, an estimator for CATE is obtained by subtracting them:
    $\hat{\tau}(X)=\hat{\mu}_{1}(X)-\hat{\mu}_{0}(X)$
  • X learner: This approach builds on the foundations of the T Learner, and starts similarly by estimating $\hat{\mu_{0}}(X)$ and $\hat{\mu_{1}}(X)$. it then uses these estimators to impute the missing potential outcomes and generate ”pseudo individual effects”:
    $if\ A=0\ :\ \ D^{0}:=\hat{\mu}_{1}(X)-Y,$
    $if\ A=1\ :\ \ D^{1}:=Y-\hat{\mu}_{0}(X)$
    Next, $D^{0}$ and $D^{1}$ are used to train two separate estimators for CATE - $\hat{\tau}_{1}(X)=\hat{\operatorname{\mathbb{E}}}[D^{1}|X]$ and $\hat{\tau}_{0}(X)=\hat{\operatorname{\mathbb{E}}}[D^{0}|X]$. Finally, a weighted average of the two estimates is used to estimate CATE. a common weighting choice is the propensity, which is also estimated from the data:
    $\hat{\tau}(X)=\hat{\pi}(X)\cdot\hat{\tau}_{0}(X)+(1-\hat{\pi}(X))\cdot\hat{\tau}_{1}(X).$
  • DR learner: This approach constructs a doubly-robust pseudo-outcome for CATE using a sub-sample of the training data, and uses the rest of the train-set to regress this pseudo outcome on X. First, using the first subset $S_{1}$ to train $\hat{\pi}(X),\ \hat{\mu}_{0}(X),\ \hat{\mu}_{1}(X)$ estimators for $\pi(X),\ \operatorname{\mathbb{E}}[Y|A=0,X],\ \operatorname{\mathbb{E}}[Y|A=1,X]$, respectively. Then, construct the following pseudo-outcome:
    $\hat{\varphi}(X,Y,A)=\frac{A-\hat{\pi}(X)}{\hat{\pi}(X)\cdot(1-\hat{\pi}(X))}(Y-\hat{\mu}_{A}(X))+\hat{\mu}_{1}(X)-\hat{\mu}_{0}(X).$
    Finally, use the rest of the train-sample, $S_{2}=S\setminus S_{1}$ to regress:
    $\hat{\tau}_{dr}(x)=\hat{\operatorname{\mathbb{E}}}[\hat{\varphi}(X,Y,A)|X=x].$
Each of the aforementioned methods can use any ML model as a ”base estimator” for any of the nuisance functions it estimates. The choice of base estimators should account for the assumed complexity of the underlying data-generating process, the size of available training data etc. Clinical trials often test complex mechanisms, which favor more flexible ML models. However, the typical sample sizes in this setting are usually quite limited, favoring simpler more regularized models. Therefore, a set of base-learners spanning a wide range of complexity levels was compared:
  • GLMs: Regularized (Lasso and Elastic-Net) linear regression was used to estimate outcome and pseudo outcomes, and logistic regression to estimate the propensity.
  • Accurate GLM (AGLM): This method utilizes Lasso regression to fit a piece-wise constant function, by first encoding each variable into nested bins [20].
  • Boosted Regression Trees: A sum of trees, where each tree is fitted on the residuals of the previous one [21].
  • Random Forests (RF): An average of trees constructed with stochastic sampling of features and sample to induce variability in the trees, which acts as a form of regularization [22].
  • Bayesian Additive Regression Trees (BART): Inspired by ensemble methods, with boosting in particular, BART also trains a sum-of-trees model, with the addition of a Bayesian regularization prior which controls the parameters of that model [23, 24].
A subset of the tested configurations are presented in the results section. Results for additional models are available in B.
Intermediate analysis of the performance of different estimators in different scenarios, revealed that no single estimator performs well across all scenarios (compared to others), with each single estimator under-performing in at least one of the scenario types. This motivated the search for an ensemble method that could provide a more robust performance than any of the methods comprising it. We used two ensemble approaches - Stacked X-Learner which uses stacking to ensemble difference prediction models within a Meta-Learner, and Consensus Based Averaging (CBA) which uses a combination of CATE estimators with high agreement.
  • Outcome Ensemble: We used stacking for the first level ($\mu_{0}$, $\mu_{1}$) and second level ($\tau_{0}$, $\tau_{1}$) of nuisance functions. The estimators used for stacking were elasticnet, RF, boosting and AGLM, which were all cross-fitted using 3-fold CV and then used as input for linear regression for the outcome on all of the training set.
  • Consensus Based Averaging (CBA): This method prioritizes models with higher agreement. Initially, predictions from each model are collected. Kendall’s Tau correlation coefficients between each pair of models are calculated, and the mean correlation coefficient for each model is determined. Models are then sorted based on their mean correlation values. The method identifies a subset of models that are in higher agreement by detecting the ”knee” in the sorted mean correlation curve. Models up to this knee point are selected, and their predictions are averaged to produce the final ensemble prediction, enhancing robustness and accuracy. For further details on the method, see appendix A.
4 Simulation Study Results
The simulation study included generating multiple synthetic scenarios with varying characteristics and levels of complexity. For each DGP, training sets of multiple sizes were randomized and all of the CATE estimation methods described in 3 were fit on each of the training sets, and then used to predict CATE on 5000 held-out evaluation patients. Each combination of scenario and training-set size was used to generate 50 random training datasets. The results reported are based on the evaluation set.
The study results demonstrate that there is no single model that is dominant across scenarios. Some models are in perfect conjunction with the DGP of some of the scenarios and therefore have an inherent advantage. For instance, the linear model-based T-learner shows dominance in larger linear scenarios, as can be seen in Figures 1 and 2.
Figure 1 shows the advantage of ensemble methods (CBA and Stacked-X Learner) in achieving competitive accuracy, achieving best or near-best performance in most tested scenarios. Figure 2 confirms this for the RoD as well.
Figure 1: Relative Access Median sRMSE
median sRMSE (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median sRMSE, compared to the best median sRMSE for that scenario.
Figure 2: Relative Access Median RoD
median RoD (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median RoD, compared to the best median RoD for that scenario.
4.1 Subgroup Evaluation
Comparing the accuracy of aggregated CATE estimation in potential subgroups supports the above findings, with the ensemble methods having good accuracy in most scenarios. Notably, in this aggregated perspective, the performance of Causal Forest are also very robust and stable, possibly due to the ”tree-like” structure of SGs used for evaluation.
Due to the overlap in patients between the potential subgroups as they were defined in section 3.2.1, the estimation errors for subgroups are not independent of each other. Nevertheless, this perspective gives intuition for the expected usefulness of an estimator for subgroup identification and evaluation.
Figure 3: Relative Access Median sRMSE on Subgroups
median sRMSE on subgroups (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median sRMSE, compared to the best median sRMSE for that scenario.
Figure 4: Relative Access Median RoD on Subgroups
median RoD (in 50 simulation runs) was calculated for different scenarios and models. For each scenario (row) the best model was identified, and the values in the figure depict the relative access median RoD, compared to the best median RoD for that scenario.
4.2 Discussion
The results of this study highlight the robustness of ensemble methods in CATE estimation, when predicting effects at both the individual patient level and in potential subgroup evaluations. We propose ensembles of two different approaches—a meta-learner with model-stacking as a base-learner, and a Consensus-Based Average of estimators, which identifies a subset of estimators with the highest agreement. Both ensemble approaches demonstrate superior performance over single models, across a wide range of scenarios.
Ensemble methods consistently achieve high accuracy and exhibit robust performance across diverse scenarios. They maintain competitive performance relative to the best individual models, even when the underlying data generating process is unknown.
Given the inherent variability in model performance and the unknown nature of the true DGP in practical settings, the robustness of ensemble methods is particularly valuable. The proposed ensembles provide reliable and accurate CATE estimations, maintaining consistent performance across diverse scenarios. This robustness is crucial for practical applications in clinical trials and personalized medicine, where accurate and stable treatment effect predictions are essential for effective decision-making.
Additional work should expand the discussion to other types of endpoints, namely binary and time-to-event endpoints, which are common in clinical research. Other ensembling approaches have been proposed in the literature for CATE estimation [26, 27, 29, 28], however their performance was only demonstrated on much larger datasets than the ones considered here. Future work will benchmark selected methods on a unified set of scenarios.
Appendix A Consensus Based Average
The proposed ensemble method combines predictions from multiple models by prioritizing those that exhibit higher agreement. Kendall’s Tau correlation coefficients are then calculated between the predictions of each pair of models, and the mean correlation coefficient for each model is determined.
Let $\mathbf{P}=[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{M}]$ be the matrix of predictions, where $\mathbf{p}_{i}$ represents the predictions from the $i$-th model, and $M$ is the number of models. Kendall’s Tau correlation coefficient between models $i$ and $j$ is given by:
$\tau_{ij}=\mathrm{Kendall^{\prime}s\;Tau}(p_{i},p_{j}).$
The mean correlation coefficient for each model $i$ is:
$\bar{\tau}_{i}=\frac{1}{M-1}\sum_{j\neq i}\tau_{ij}.$
Next, the models are sorted in descending order based on their mean correlation values. To identify a subset of models that are in higher agreement, the method detects the ”knee” in the sorted mean correlation curve, which is the point where the difference between consecutive mean correlation values is minimized. Let $\bar{\tau}$ be the sorted mean correlation values:
$\bar{\tau}=[\bar{\tau}_{(1)},\bar{\tau}_{(2)},...\bar{\tau}_{(M)}].$
The ”knee” is detected by finding the index $k$ where the difference between consecutive mean correlation values is minimized:
[…]
models up to this ”knee” point are selected for the final ensemble. The predictions from these selected models, which have higher mean correlations and thus more agreement, are averaged to produce the final ensemble prediction. Let $\mathcal{S}$ be the set of selected models up to the knee point:
$\mathcal{S}=\{p_{(1)},p_{(2)},\ldots,p_{(k)}\}.$
The final ensemble prediction $\mathbf{p}_{\text{ensemble}}$ is:
$CBA=\frac{1}{k}\sum_{p_{i}\in S}p_{i}.$
By averaging the predictions from these selected models, the ensemble method leverages the consensus among models, enhancing the robustness and accuracy of the ensemble’s performance.
Appendix B Supplementary Results
Scenario p n CF-DML CF-DML S-BART T-Linear X-AGLM X-BART X-Boosting X-RF DR-RF Stacked-X CBA
PD-L1 5 100 0.88 0.79* 1.20 0.80 0.94 1.03 0.92 0.97 1.19 0.81 0.81
250 0.69 0.62* 0.98 0.73 0.80 0.75 0.68 0.74 0.81 0.65 0.66
500 0.57 0.45* 0.93 0.72 0.84 0.65 0.55 0.65 0.69 0.55 0.58
Highly NL 10 100 1.04 1.18 1.01* 1.19 1.11 1.08 1.34 1.05 1.21 1.11 1.05
250 0.98 1.02 1.01 1.16 0.93 0.93* 1.14 0.94 1.02 0.97 0.93
500 0.90 0.88 1.01 1.08 0.77* 0.82 0.94 0.85 0.91 0.78 0.84
750 0.84 0.82 1.00 1.05 0.67 0.78 0.82 0.79 0.83 0.65* 0.78
Slightly NL 10 100 0.88 0.76 0.99 0.85 0.80 0.72 0.96 0.79 0.86 0.87 0.67*
250 0.61 0.58 0.95 0.63 0.57 0.51* 0.72 0.55 0.61 0.54 0.52
500 0.44 0.50 0.94 0.55 0.44 0.42 0.58 0.46 0.50 0.39* 0.42
Slightly NL 20 100 0.98 0.92 0.99 1.02 0.90 0.77* 0.93 0.85 1.51 0.89 0.78
250 0.76 0.62 0.98 0.71 0.61 0.55* 0.72 0.72 0.71 0.58 0.59
500 0.55 0.50 0.97 0.65 0.50 0.47* 0.60 0.62 0.59 0.48 0.48
Linear 10 100 0.99 1.07 1.00 1.00 1.04 0.94 1.34 0.92 1.13 1.09 0.91*
250 0.86 0.85 0.98 0.59* 0.85 0.73 1.06 0.81 0.84 0.67 0.73
500 0.76 0.74 0.97 0.41* 0.67 0.59 0.85 0.71 0.74 0.46 0.60
Table 1: median sRMSE (over 50 simluated datasets)
Scenario p n CF-DML CF-DML S-BART T-Linear X-AGLM X-BART X-Boosting X-RF DR-RF Stacked-X CBA
PD-L1 5 100 0.25* 0.25 0.40 0.28 0.33 0.32 0.34 0.35 0.38 0.29 0.28
250 0.23 0.23* 0.32 0.25 0.28 0.28 0.26 0.28 0.28 0.25 0.24
500 0.22 0.21* 0.30 0.24 0.33 0.25 0.23 0.24 0.25 0.22 0.22
Highly NL 10 100 0.45 0.44 0.48 0.50 0.44 0.41 0.41* 0.43 0.42 0.46 0.42
250 0.38 0.36 0.47 0.50 0.34* 0.35 0.36 0.36 0.37 0.36 0.35
500 0.34 0.32 0.47 0.50 0.27* 0.29 0.31 0.31 0.33 0.27 0.30
750 0.30 0.29 0.47 0.50 0.24 0.27 0.28 0.28 0.29 0.23* 0.27
Slightly NL 10 100 0.21 0.24 0.38 0.23 0.25 0.22 0.27 0.23 0.26 0.25 0.20*
250 0.15 0.21 0.32 0.18 0.19 0.16 0.23 0.17 0.20 0.15 0.15*
500 0.14 0.19 0.32 0.16 0.15 0.14 0.20 0.16 0.17 0.13* 0.14
Slightly NL 20 100 0.33 0.31 0.41 0.29 0.28 0.25 0.28 0.27 0.34 0.26 0.25*
250 0.19 0.22 0.39 0.22 0.21 0.18 0.24 0.22 0.25 0.18 0.18*
500 0.16 0.19 0.39 0.20 0.17 0.15 0.20 0.18 0.19 0.15 0.15*
Linear 10 100 0.35 0.33 0.42 0.32 0.37 0.30* 0.34 0.33 0.35 0.36 0.32
250 0.26 0.27 0.39 0.16* 0.28 0.23 0.29 0.26 0.28 0.20 0.22
500 0.23 0.23 0.39 0.11* 0.21 0.18 0.25 0.23 0.24 0.13 0.17
Table 2: median RoD (over 50 simluated datasets)
Appendix C PD-L1-biological data generation model
Figure 5 illustrates the PD-L1 (Programmed Death Ligand 1) model, which delineates prominent factors influencing tumor progression. This model integrates various biological and immunological elements to simulate some key aspects of the dynamics of tumor growth.
  • Mutation Burden: One of two ”root” elements in our model, this factor impacts both the post treatment CD8+ levels, as well as potentially impacting Tumor Growth in additional pathways not modeled here, thus represented as a direct impact on Tumor Growth.
  • Immune Phenotype: The other ”root” element, this ordinal variable classifies the presence of immune cells in the tumor area into 3 classes - ”inflamed”, ”excluded” or ”desert” - indicating decreasing levels of immune activity. Higher levels of activity act to increase TGF-$\beta$ (Transforming Growth Factor-beta), basleine CD8+ levels and PD-L1 levels.
  • TGF-$\mathbf{\beta}$: Causes the CD8+ levels to increase, regardless of the treatment.
  • CD8+ Effector: We model the CD8+ levels post-treatment as the baseline levels, with the addition of terms relating to TGF-$\beta$ and to an interaction between mutation burden, and the change in PD-L1 caused by the treatment.
  • PD-L1: We model the ligand levels as an ordinal factor with 3 levels. The model assumes a 1 level decrease with treatment.
  • Tumor Growth: The tumor growth is driven upward by the mutation burden, and downward by the post-treatment levels of CD8+ effector.
[…]
Figure 5: PD-L1 Model, relating some key factors for tumor progression
Shaded nodes compose the features $X$. Tumor Growth (thick circle on the right) is the endpoint $y$.